# Precise Robotic Weed Spot-Spraying for Reduced Herbicide Usage and Improved Environmental Outcomes - A Real-World Case Study

### v1

@@ 1 Introduction
Herbicides are a serious threat to non-target plants and animals because they are readily carried by water run-off from farmland into rivers, creeks, coastal and inshore areas. Consequently, herbicide usage has attracted significant public attention and pressure from regulators Mohring2020 . A promising approach to reduce herbicide runoff is to use precision agriculture and digital technologies Allmendinger2022 ; Tataridas2022 ; Wang2020 . Examples of such technologies include the use of Unoccupied Aerial Vehicles (UAV) for precise weed mapping Zhang2021 , non-chemical robotic weed management VanDerWeide2008 , and the use of UAVs Rai2022 and ground vehicles Calvert2021 for spot-spraying of weeds. Spot spraying allows for a much more efficient use of chemicals since herbicides can be sprayed only where required, and not onto the bare ground or non-target plants.
@@
In many scenarios, herbicides must be applied during the growth of the crop. This requires the spot sprayer to discriminate between weed and crop plants, a situation which is called green-on-green detection. This differs from so-called green-on-brown detection, which targets all plants indiscriminately. For instance, a common green-on-brown detection technology is Near-Infrared (NIR) optical sensing, which has been implemented in systems such as WeedSeeker Trimble and Weed-IT WeedIT . These systems react to the NIR light emitted by plants during photosynthesis. As a result, these systems do not discriminate between plant types and therefore cannot separate crop from weed. Hence, they can only be applied in a green-on-brown scenario, such as in fallow paddocks.
@@
Other sensor technologies for weed detection include multi-spectral, hyper-spectral, 3D, and depth cameras, as well as stereo and ultrasonic sensors Gerhards2022 ; Esposito2021 . The most commonplace sensor that can be used to detect weeds in a green-on-green scenario is a standard Red-Green-Blue (RGB) camera, which works in the visible light spectrum Esposito2021 . Precise detection of weeds among surrounding green plants, which can have similar features, is a complex problem in Computer Vision (CV). Fortunately, with the advent of Deep Learning (DL) technology and Convolutional Neural Networks (CNN) in the past decade, weed detection with image processing has progressed significantly Yu2019 ; Lammie2019 ; Olsen2019 ; Calvert2021 ; Allmendinger2022 .
@@
In this paper, we describe the development of a ground-based spot-spraying tool for sugarcane, which we call AutoWeed. Our system is illustrated in Fig. 1. We developed the computer vision algorithms for weed detection and designed the spraying system to be retrofitted to existing farm machinery. We carried out field trails in the Burdekin region of Queensland, Australia, which is one of Australia’s largest sugarcane growing regions. We conducted a comprehensive set of experiments to measure the real-world impact of spot spraying in terms of the reduction of herbicide used, the resulting weed knockdown efficacy, and the measured concentration of herbicide active ingredients in water runoff induced by irrigation.
@@ 2.1 Deep learning and its application to precision agriculture
Deep learning has grown to be a prevalent technology that has brought astonishing performance to computer vision. Deep learning-enabled CV algorithms are currently being applied in various domains ranging from aquaculture Saleh2022 to marine Laradji2021 to environmental Jahanbakht2022 , and to medical sciences Azghadi2020 ; Mustafa2021 . Similarly, in the past few years, computer vision for precision agriculture and weed detection has significantly grown. For instance, Lu et al. have conducted a survey of public datasets for computer vision tasks in precision agriculture using machine and deep learning Lu2020 . They have surveyed 10 studies for fruit detection, disease, damage and flower detection and counting, as well as yield prediction, using deep learning.
@@
Among the precision agriculture applications that Lu et al. have surveyed, weed control has emerged as a major priority. They reported that there have been 15 publicly available ground and aerial image datasets collected in field using various sensors, such as RGB cameras, multi-spectral and multi-modal sensors, in the past 8 years. Several of these datasets such as Haug2014 ; Lameski2017 ; Chebrolu2017 ; SantosFerreira2017 ; DiCicco2017 ; Sa2017 ; Lottes2018 ; Teimouri2018 ; Sa2018 ; Olsen2019 ; EspejoGarcia2020 have been used in deep learning-based weed control tasks McCool2017 ; Bosilj2020 ; SantosFerreira2019 ; DiCicco2017 ; Sa2017 ; Lottes2020 ; Teimouri2018 ; Sa2018 ; Olsen2019 ; EspejoGarcia2020 . Here, weed control can be any use of computer vision and deep learning to segment McCool2017 ; Lameski2017 ; Bosilj2020 , or simply classify Olsen2019 ; SantosFerreira2017 weeds for mapping Sa2018 or spraying Calvert2021 .
@@
Other studies have demonstrated the potential of deep learning for weed detection and control in precision agriculture. For instance, Jin et al. Jin2022 developed a deep learning-based method for detecting herbicide weed control spectrum in turfgrass, which can be used in a machine vision-based autonomous spot-spraying system of smart sprayers. Similarly, Harders et al. harders2022deep proposed a deep learning approach for UAV-based weed detection in horticulture using edge processing and presented experimental results. In addition to detecting weeds, researchers have also explored the use of deep learning to generate synthetic weed images. Chen et al. chen2023deep applied diffusion probabilistic models to generate high-quality synthetic weed images based on transfer learning, while Divyanth et al. divyanth2022image aimed to reduce the effort needed to prepare large image datasets by creating artificial images of maize and common weeds through conditional Generative Adversarial Networks. Other researchers have focused on developing autonomous systems for weed detection and control. Patel et al. patel2022design described the design of an autonomous agricultural robot for real-time weed detection using CNN, while Gao et al. gao2022transferring developed a deep convolutional network that enables the prediction of both field and aerial images from UAVs for weed segmentation and mapping with only field images provided in the training phase. Narayana et al. narayana2023efficient also developed an efficient real-time weed detection technique using deep learning. To facilitate the development of deep learning-based weed detection methods, Wang et al. wang2022weed25 presented Weed25, a deep learning dataset for weed identification containing 14,035 images of 25 different weed species. Finally, Murad et al. murad2023weed conducted a systematic literature review on current state-of-the-art DL techniques for weed detection.
@@
In this paper, we use weed classification for spot-spraying. Classification, as opposed to segmentation, is an easier computer vision task that can be carried out very accurately using deep learning Olsen2019 ; Calvert2021 . The goal here is to identify if an image includes one or more weed(s) of interest.
@@ 2.2 DL for weed spot-spraying
Although many works have investigated and reported the potential of using CNNs for weed classification in different scenarios within on-ground Yu2019 ; Jin2022 ; Zhang2018 ; Knoll2019 ; Subeesh2022 or UAV images De2021 ; Huang2020 ; Beeharry2020 ; Khan2021 ; Rai2022 , research works on in-field weed detection in conjunction with spot-spraying has not seen wide exploration. This is mainly due to the significant efforts required to perform efficient in-field trials outside a lab environment. Examples of ground systems that have used CNNs for in-field weed spot-spraying include the following.
@@
In one of the earliest studies Partel2019 , Partel et al. developed a prototype spot-spraying system and simulated it using two scenarios. In the first scenario, a vegetable field was simulated using artificial weeds (targets) and artificial plants (non-targets), while in the second scenario, they applied their prototype to real plants. They also investigated the weed detection performance and system operations using two different Graphics Processing Units (GPUs) reporting the achieved accuracies. Finally, they utilized a GPS device integrated into their detection system to produce weed maps. Although this showcased the use of deep learning technology to perform spot-spraying of weeds, it was in a controlled and simulated environment.
@@
In ref. Liu2021 , a variable rate chemical spraying system was partially trialled in the field at a low application speed of 3 km/hr and using a prototype system controlling chemical usage in a strawberry crop. However, this was only a prototype design that did not report herbicide reduction achieved or the potential water quality benefits that could be obtained.
@@
In ref. Du2022 , a CNN-based multi-class under-canopy weed control robotic unit named SAMBot was developed and trialled in a limited setting in flax fields with medium weed density. SAMBot was able to achieve an average weed classification accuracy of 90%. It also showed reduced herbicide usage compared to a commercial sprayer. This study trialled the robot in only 15 m of fields and further results should be shown before the efficacy of the system could be fully demonstrated.
@@
In a recent study, Allmendinger et al. reviewed different commercial technologies and prototypes for precision patch spraying and spot spraying Allmendinger2022 . They demonstrated that weed spot-spraying using DL has seen a lot of attention from the industry and several systems are currently being trialled and made commercially available. Examples of these products that use RGB cameras and DL include Bilberry, Blue River’s See and Spray, EcoPatch, and Greeneye. For a detailed review of these systems, please see Allmendinger2022 .
@@ 3.1 The AutoWeed system
Figure 2: The AutoWeed weed detection and spraying boom mount design for a 1-metre boom section, including (a) the AutoWeed detection unit, (b) a protective sun shade, (c) a 1“ wet boom, (d) a 40 $\times$ 40 mm steel hollow section frame, and (e) TeeJet solenoids and nozzle body adaptors.
@@
These units were purpose-built for retrofitting to existing agricultural vehicles. Each unit comprises a machine vision camera, an NVIDIA Jetson embedded GPU processor, and a custom solenoid sprayer board that can individually control up to four solenoids per camera. The design is enclosed in a robust PVC housing attached to a mounting plate to protect the internals from the harsh environment. The units can be mounted to any 40 mm or 50 mm steel hollow section frame and is compatible with the commercially-available TeeJet boom and spray nozzle components. For weed detection, the units utilise deep learning based image classification models that are trained to weeds of interest using large labelled image datasets, following our previously reported methodology Olsen2019 .
@@
In this project, the AutoWeed detection and spraying prototypes were retrofitted for broadcast spraying and Irvin leg spraying on two spraying vehicles. The first was a large 13-row high-rise John Deere R4720 self-propelled sprayer (see Figure 3). The second was a smaller 4-row three-point linkage boom sprayer towed behind a Landini tractor (as shown in Figure 4). For both systems, the AutoWeed detection units were mounted in front of the sprayers and positioned between the sugarcane row centres.
@@
Figure 3: (a) A 13-row high-rise John Deere R4720 self-propelled sprayer retrofitted with AutoWeed detection prototypes for Irvin leg spraying in sugarcane. (b) The AutoWeed detection units were mounted directly in front of the Irvin legs and positioned between the cane row centres.
Figure 4: A 4-row three-point linkage boom sprayer retrofitted with AutoWeed detection prototypes for weed spot-spraying in sugarcane. The AutoWeed detection units were mounted to the spray boom frame, which can be utilised for Irvin leg or broadcast spraying
3.2 The weed dataset
@@
One of the crucial first steps when developing a deep learning algorithm is collecting suitable datasets. The quantity, quality, and diversity of the data used to train the model directly impacts its performance. For weed detection and spot-spraying, the collected dataset should contain diverse images of both crops and weeds in different growth stages, lighting conditions, and environments. The larger and more diverse the dataset, the better the model will be able to generalise and accurately identify weeds in new, unseen situations. It is also important to carefully label the data to ensure that the model is properly trained on the correct classifications Olsen2019 . In spot-spraying applications using deep learning, the dataset is usually collected from a specific site, for which a specific model is trained. This method is commonly known as site-specific weed control and is effective because it takes into account specific conditions of a site including the growth stage and variety in shape and features of the same weeds, which may be starkly different in two different sites.
@@
For the project presented in this paper, we collected a total of 1,447,456 site-specific images before each of the respective six spray trials as described in Table 1. This data allowed training site-specific deep learning algorithms to detect target weeds for the different sugarcane paddocks. Data collection was performed in the Burdekin region of Queensland, Australia using an All-Terrain Vehicle (ATV) retrofitted with AutoWeed detection units, similar to that shown in Calvert2021 . For adaptation to sugarcane, the AutoWeed detection units were retrofitted onto spray booms centred on the interrow (i.e. in between crop rows). This allowed clear vision of weeds in the row underneath the sugarcane leaf canopy and in the interrow soil area, as shown in Fig. 5. A tiled classification approach was chosen whereby images were split into 2 $\times$ 2 tiles for annotation and 1 $\times$ 2 tiles for inference. This allowed rapid collection and annotation of high tile counts during dataset collection, and high-speed performance with a batch size of two during inference.

### v2

# Precision Robotic Spot-Spraying: Reducing Herbicide Use and Enhancing Environmental Outcomes in Sugarcane

@@ 1 Introduction
Herbicides are a serious threat to non-target plants and animals because they are readily carried by water run-off from farmland into rivers, creeks, coastal and inshore areas. Consequently, herbicide usage has attracted significant public attention and pressure from regulators. A promising approach to reduce herbicide runoff is to use precision agriculture and digital technologies [1]. Examples of such technologies include the use of Unoccupied Aerial Vehicles (UAV) for precise weed mapping, non-chemical robotic weed management, and the use of UAVs [2] and ground vehicles [3] for spot-spraying of weeds. Spot spraying allows for a much more efficient use of chemicals since herbicides can be sprayed only where required, and not onto the bare ground or non-target plants.
@@
In many scenarios, herbicides must be applied during the growth of the crop. This requires the spot sprayer to discriminate between weed and crop plants, a situation which is called green-on-green detection. This differs from so-called green-on-brown detection, which targets all plants indiscriminately. For instance, a common green-on-brown detection technology is Near-Infrared (NIR) optical sensing, which has been implemented in systems such as WeedSeeker [4] and Weed-IT [5]. These systems react to the NIR light emitted by plants during photosynthesis. As a result, these systems do not discriminate between plant types and therefore cannot separate crop from weed. Hence, they can only be applied in a green-on-brown scenario, such as in fallow paddocks.
@@
Other sensor technologies for weed detection include multi-spectral, hyper-spectral, 3D, and depth cameras, as well as stereo and ultrasonic sensors. The most commonplace sensor that can be used to detect weeds in a green-on-green scenario is a standard Red-Green-Blue (RGB) camera, which works in the visible light spectrum. Precise detection of weeds among surrounding green plants, which can have similar features, is a complex problem in Computer Vision (CV). Fortunately, with the advent of Deep Learning (DL) technology and Convolutional Neural Networks (CNN) in the past decade, weed detection with image processing has progressed significantly [6].
@@
In this paper, we describe the development of a ground-based spot-spraying tool for sugarcane, which we call AutoWeed. Our system is illustrated in Fig. 1. We developed the computer vision algorithms for weed detection and designed the spraying system to be retrofitted to existing farm machinery. We carried out field trials in the Burdekin region of Queensland, Australia, which is one of Australia’s largest sugarcane growing regions.
We conducted a comprehensive set of experiments to measure the real-world impact of spot spraying in terms of the reduction of herbicide used, the resulting weed knockdown efficacy, and the measured concentration of herbicide active ingredients in water runoff induced by irrigation. We hypothesised that:
  1. 1.
    Spot spraying is more efficient in terms of herbicide usage compared to blanket spraying.
  2. 2.
    Spot spraying provides comparable weed knockdown efficacy compared to blanket spraying.
  3. 3.
    Spot spraying results in lower concentrations of herbicide active ingredients in water runoff compared to blanket spraying.
  4. 4.
    Spot spraying positively impacts crop yield compared to traditional blanket spraying methods.
By addressing these hypotheses, we aim to provide a comprehensive evaluation of the efficacy and efficiency of the AutoWeed system.
@@ 2.1 Deep learning and its application to precision agriculture
Deep learning has grown to be a prevalent technology that has brought astonishing performance to computer vision. Deep learning-enabled CV algorithms are currently being applied in various precision agriculture domains.
@@
Similarly, in the past few years, computer vision for precision agriculture and weed detection has significantly grown. For instance, Lu et al. have conducted a survey of public datasets for computer vision tasks in precision agriculture using machine and deep learning [7]. They have surveyed 10 studies for fruit detection, disease, damage and flower detection and counting, as well as yield prediction, using deep learning.
@@
Among the precision agriculture applications that Lu et al. have surveyed, weed control has emerged as a major priority. They reported that there have been 15 publicly available ground and aerial image datasets collected in field using various sensors, such as RGB cameras, multi-spectral and multi-modal sensors, in the past 8 years. Several of these datasets such as [8] have been used in deep learning-based weed control tasks [6]. Here, weed control can be any use of computer vision and deep learning to segment, or simply classify weeds for mapping or spraying.
@@
Other studies have demonstrated the potential of deep learning for weed detection and control in precision agriculture. For instance, Jin et al. [9] developed a deep learning-based method for detecting herbicide weed control spectrum in turfgrass, which can be used in a machine vision-based autonomous spot-spraying system of smart sprayers. Similarly, Harders et al. [10] proposed a deep learning approach for UAV-based weed detection in horticulture using edge processing and presented experimental results. In addition to detecting weeds, researchers have also explored the use of deep learning to generate synthetic weed images. Chen et al. [11] applied diffusion probabilistic models to generate high-quality synthetic weed images based on transfer learning, while Divyanth et al. [12] aimed to reduce the effort needed to prepare large image datasets by creating artificial images of maize and common weeds through conditional Generative Adversarial Networks. Other researchers have focused on developing autonomous systems for weed detection and control. Patel et al. [13] described the design of an autonomous agricultural robot for real-time weed detection using CNN, while Gao et al. [14] developed a deep convolutional network that enables the prediction of both field and aerial images from UAVs for weed segmentation and mapping with only field images provided in the training phase. Narayana et al. [15] also developed an efficient real-time weed detection technique using deep learning. To facilitate the development of deep learning-based weed detection methods, Wang et al. [16] presented Weed25, a deep learning dataset for weed identification containing 14,035 images of 25 different weed species. Finally, Murad et al. [17] conducted a systematic literature review on current state-of-the-art DL techniques for weed detection.
@@
In this paper, we use weed classification for spot-spraying. Classification, as opposed to segmentation, is an easier computer vision task that can be carried out very accurately using deep learning [6, 3]. The goal here is to identify if an image includes one or more weed(s) of interest.
@@ 2.2 DL for weed spot-spraying
Although many works have investigated and reported the potential of using CNNs for weed classification in different scenarios within on-ground [18] or UAV images, research works on in-field weed detection in conjunction with spot-spraying has not seen wide exploration. This is mainly due to the significant efforts required to perform efficient in-field trials outside a lab environment. However, recent advancements in deep learning have shown promising results in the detection and classification of weeds, which is crucial for the development of targeted spraying systems.
@@
In one of the earliest studies [19], Partel et al. developed a prototype spot-spraying system and simulated it using two scenarios. In the first scenario, a vegetable field was simulated using artificial weeds (targets) and artificial plants (non-targets), while in the second scenario, they applied their prototype to real plants. They also investigated the weed detection performance and system operations using two different Graphics Processing Units (GPUs) reporting the achieved accuracies. Finally, they utilized a GPS device integrated into their detection system to produce weed maps. Although this showcased the use of deep learning technology to perform spot-spraying of weeds, it was in a controlled and simulated environment.
@@
In ref. [20], a variable rate chemical spraying system was partially trialled in the field at a low application speed of 3 km/hr using a prototype system controlling chemical usage in a strawberry crop. However, this was only a prototype design that did not report herbicide reduction achieved or the potential water quality benefits that could be obtained.
@@
In ref. [21], a CNN-based multi-class under-canopy weed control robotic unit named SAMBot was developed and trialled in a limited setting in flax fields with medium weed density. SAMBot was able to achieve an average weed classification accuracy of 90%. It also showed reduced herbicide usage compared to a commercial sprayer. This study trialled the robot in only 15 m of fields and further results should be shown before the efficacy of the system could be fully demonstrated.
While significant advancements have been made in the application of deep learning for weed detection, the integration of these technologies into in-field weed spot-spraying systems remains under-explored. Future research should focus on conducting more extensive field trials and reporting on the practical benefits such as herbicide reduction and water quality improvements, which are the objectives of our paper.
2.3 Current Status of Research on Target Spraying Robots
@@
Target spraying robots have become a pivotal technology in precision agriculture, significantly improving the accuracy and efficiency of herbicide application. In a recent study, Allmendinger et al. reviewed different commercial technologies and prototypes for precision patch spraying and spot spraying [1]. They demonstrated that weed spot-spraying using DL has seen a lot of attention from the industry and several systems are currently being trialled and made commercially available. For a detailed review of these systems, please see [1]. Key developments in target spraying robots include:
Autonomous Navigation and Detection Systems: Modern target spraying robots are equipped with advanced navigation systems and real-time weed detection capabilities. For instance, autonomous robots like the one developed by [22] use GPS and LiDAR to navigate fields and identify weed locations accurately.
Integration of Deep Learning Models: The use of deep learning models, such as CNNs, for weed detection has been a game-changer. Systems like those reviewed by [23] utilize DL models to distinguish between crop and weed species with high precision, enabling targeted herbicide application.
Field Trials and Practical Implementations: Extensive field trials have been conducted to test the efficacy of target spraying robots under real-world conditions. Studies by [24] demonstrated significant herbicide savings and improved weed control efficacy in large-scale agricultural settings.
Commercialization and Industry Adoption: Several commercial products have emerged from this research, including Bilberry [25], Blue River’s See and Spray [26], and EcoPatch [27]. These systems are being adopted by farmers globally, driven by their potential to reduce chemical usage and enhance crop yields.
@@ 3.1 The AutoWeed system
Figure 2: The AutoWeed weed detection and spraying boom mount design for a 1-metre boom section, including (a) the AutoWeed detection unit including a machine vision camera, an NVIDIA Jetson GPU, and a custom solenoid sprayer board that can individually control up to four solenoids per camera; (b) a protective sun shade, (c) a 1“ wet boom, (d) a 40 $\times$ 40 mm steel hollow section frame, and (e) TeeJet solenoids and nozzle body adaptors.
@@
These units were purpose-built for retrofitting to existing agricultural vehicles. Each unit comprises a machine vision camera, an NVIDIA Jetson embedded GPU processor, and a custom solenoid sprayer board that can individually control up to four solenoids per camera. The design is enclosed in a robust PVC housing attached to a mounting plate to protect the internals from the harsh environment. The units can be mounted to any 40 mm or 50 mm steel hollow section frame and are compatible with the commercially-available TeeJet boom and spray nozzle components.
@@
In this project, the AutoWeed detection and spraying prototypes were retrofitted for broadcast spraying and Irvin leg spraying on two spraying vehicles. The first was a large 13-row high-rise John Deere R4720 self-propelled sprayer (see Figure 3, a and b). The second was a smaller 4-row three-point linkage boom sprayer towed behind a Landini tractor (as shown in Figure 3, c and d). For both systems, the AutoWeed detection units were mounted in front of the sprayers and positioned between the sugarcane row centres.
3.2 Weed image acquisition, analysis, and application technology
The AutoWeed units utilise a machine vision camera pointing directly down to capture images of the field (see Fig. 2). The height of the camera in the detection unit, as shown in Fig. 1(b), is set based on the required field of view of the target crop row width. For sugarcane paddocks in the Burdekin region, the row width usually varies between 1.5 to 1.6 metres. To achieve a field of view that captured the full sugarcane row width, the camera lens was positioned at least 1m above the ground to ensure a horizontal field of view of at least 1.6 metres. We also utilised separately adjustable mounting for the camera and the spray nozzles so that the height of the spray nozzle to the target weeds could be adjusted without affecting the height of the camera.
Deep learning image classification models are trained using large labelled image datasets collected from the fields and weeds of interest. These models are then used in-field to analyse the images collected in real-time during spot-spraying and detect those containing weeds of interest to spot-spray them using our custom solenoid sprayer board that can individually control up to four solenoids per detection unit (see Fig. 2e). This process is explained in detail below.
(a) John Deere R4720 sprayer with AutoWeed
(b) Detection units on Irvin legs
(c) 4-row sprayer for broadcast spraying
(d) Detection units on spray boom frame
@@ 3.2 Weed image acquisition, analysis, and application technology
Figure 3: Illustration of the AutoWeed system retrofitted on different sprayers: (a) a 13-row high-rise John Deere R4720 self-propelled sprayer for Irvin leg spraying, with (b) detection units mounted in front of the Irvin legs between cane row centres; (c) a 4-row sprayer fitted for broadcast spraying, with (d) detection units mounted to the spray boom frame.
3.2.1 The weed dataset
@@
One of the crucial first steps when developing a deep learning algorithm is collecting suitable datasets. The quantity, quality, and diversity of the data used to train the model directly impact its performance. For weed detection and spot-spraying, the collected dataset should contain diverse images of both crops and weeds in different growth stages, lighting conditions, and environments. The larger and more diverse the dataset, the better the model will be able to generalise and accurately identify weeds in new, unseen situations. It is also important to carefully label the data to ensure that the model is properly trained on the correct classifications [6]. In spot-spraying applications using deep learning, the image dataset is typically collected from a specific site and focuses on the weed(s) of interest, for which a dedicated model is trained. This method, known as site-specific weed control, is effective because it considers the unique conditions of each site, including the growth stage and the variety in shape and features of the same weeds, which can differ significantly between locations.
@@
For the project presented in this paper, we collected a total of 1,447,456 site-specific images before each of the respective six spray trials as described in Table 1. This data allowed training site-specific deep learning algorithms to detect target weeds for the different sugarcane paddocks. Data collection was performed in the Burdekin region of Queensland, Australia using an All-Terrain Vehicle (ATV) retrofitted with AutoWeed detection units that include a machine vision camera pointing downward, similar to that shown in [3]. For adaptation to sugarcane, the AutoWeed detection units were retrofitted onto spray booms centred on the interrow (i.e. in between crop rows). This allowed clear vision of weeds in the row underneath the sugarcane leaf canopy and in the interrow soil area, as shown in Fig. 4. Each dataset collection took approximately one hour, during which a few hundred thousand images (see Table 1) were collected from several sugarcane rows.
In the six trials presented in our paper, we focused on annual grasses, broadleaf weeds and nutgrass. The weeds present in our trials were selected based on direct consultation with the participating sugarcane grower, and were common weeds of sugarcane paddocks in the Burdekin region of QLD, Australia. As shown in Table 1, nutgrass (Cyperus rotundus) has been one of the main weeds targeted in our project. This sedge can be controlled in a sugarcane crop by applying a selective and costly herbicide treatment (halosulfuron-methyl) and shares similar colour and shape features with early stage sugarcane, which makes it an interesting target for spot-spraying applications using deep learning among sugarcane crops. Other weeds of interest included two grass species: summer grass (Digitaria ciliaris) and crowsfoot (Eleusine indica); and broadleaf weeds: Red convolvulus (Ipomoea hederifolia), giant pigweed (Trianthema portulacastrum) and sesbania pea (Sesbania cannabina), which we treated in a mung bean rotation crop. It is worth noting that while each paddock may have different weeds, targeting these specific weeds was prioritised due to their higher impact and herbicide costs. Additionally, in our trial, we grouped all grass and broadleaf weeds in two respective categories, as there was no need to spot-spray specific species of broadleaf or grass weeds.